# Asymptotic stability equals exponential stability—while you twist your eyes

### v1

@@ 3   Stability preserving homotopies
Remark 3.4 (On weaker notions of stability).
We note that, in general, one cannot relax global asymptotic stability to mere stability. A reason being that the (Hopf) index of corresponding equilibria under those vector fields is by no means fixed to $(-1)^{n}$ [ref:krasnosel1984geometrical, Sec. 52] (for $n$ the dimension of the state space), that is, since this index is homotopy invariant, if the indices are different, no homotopy exists between them. This also means that these equilibria cannot just be homotoped to equilibria that are GAS. Interestingly, Krasnosel’skiĭ and Zabreĭko considered precisely that question: “We now turn to the following question: Consider an autonomous system such that zero is an equilibrium state which is only known to be Ljapunov-stable. Is it always possible to deform this system into a system such that zero is an asymptotically stable equilibrium point?” [ref:krasnosel1984geometrical, p. 342]$\circ$
Next, we briefly comment on the particular class of semiflows we consider.
Remark 3.5 (On flows generated by vector fields).
Converse Lyapunov theory for flows is intimately connected to vector fields, e.g., see [ref:fathi2019smoothing]. For instance, techniques to smooth a continuous Lyapunov function for a flow naturally result in a smooth function that allows for asserting stability under the vector field that generates this flow. Secondly, there are flows—not generated by a continuous vector field—that fail to admit a smooth Lyapunov function [ref:fathi2019smoothing, Sec. 7]. This motivated us to work with (semi)flows generated by vector fields. It is unclear if a generalization exists. Note, we are not talking about maps. $\circ$
@@ 4   A view from optimal transport
Let $0\in\mathbb{R}^{n}$ be GAS under the smooth ODE $\dot{x}=X(x)$. As such, we know that there must be a smooth Lyapunov function $V$, asserting stability of 00. We can simply construct a straight-line homotopy from $X$ to $-\nabla V$, i.e., $[0,1]\ni s\mapsto(1-s)X-s\nabla V$ and preserve stability throughout (since $V$ remains a valid Lyapunov function). Now it is tempting to believe that we should be able to homotope $x\mapsto V(x)$ into $x\mapsto V_{q}(x):=\frac{1}{2}\|x\|_{2}^{2}$ along “Lyapunov functions”, specifically, along invex functions such that their negative gradient flow is well-defined and renders 00 GAS. By passing to semiflows instead of vector fields, Theorem 3.2 confirms this belief.
However, to refine our understanding towards a similar result for vector fields, we note that our desire relates to optimal transport (OT) in the following way. Let $\mu_{0},\mu_{1}\in\mathscr{P}(\mathbb{R}^{n})$ be Borel probability measures on $\mathbb{R}^{n}$. Now suppose that $\mathrm{d}\mu_{0}=\kappa_{0}e^{-V}\mathrm{d}\lambda^{n}$ and $\mathrm{d}\mu_{1}=\kappa_{1}e^{-V_{q}}\mathrm{d}\lambda^{n}$, for $\kappa_{0},\kappa_{1}$ normalization constants, $\lambda^{n}$ the Lebesgue measure on $\mathbb{R}^{n}$ and $e^{-V},e^{-V_{q}}$ (unnormalized) densities corresponding to our Lyapunov functions. Can we transport $\mu_{0}$ to $\mu_{1}$ along sufficiently regular measures, preserving unimodality?
Suppose the answer is yes, then we can construct a path of densities $[0,1]\ni s\mapsto f(x;s)$, with $f(x;0)=e^{-V(x)}$ and $f(x;1)=e^{-V_{q}(x)}$. Now, if $f$ is sufficiently regular, we can conclude that $\nabla\log f(x;s)$ is a smooth vector field, rendering 00 GAS, for each fixed $s$. Thus, understanding the transportation of measures can provide further insights. It is particularly interesting that the standard Gaussian measure results in the canonical ODE $\dot{x}=-x$.
In what follows we touch upon this viewpoint and show how OT and our homotopy questions are indeed intimately related.
The “Monge formulation”, with a quadratic cost on $\mathbb{R}^{n}$, of OT is as follows, let $\mu,\nu\in\mathscr{P}_{2}(\mathbb{R}^{n}):=\{\mu\in\mathscr{P}(\mathbb{R}^{n}):\int\|x\|_{2}^{2}\mathrm{d}\mu(x)<+\infty\}$, then we would like to solve
$\inf_{\{T:T\#\mu=\nu\}}\int_{\mathbb{R}^{n}}\|x-T(x)\|_{2}^{2}\mathrm{d}\mu(x),$ (4.1)
where $T\#\mu$ denotes the pushforward of $\mu$, i.e., $T\#\mu(A)=\mu(T^{-1}(A))$ for all Borel sets $A\subseteq\mathbb{R}^{n}$. Now, Brenier’s seminal work [ref:brenier1991polar], showed that if $\mu\ll\lambda^{n}$, then, there is always a convex map $\phi:\mathbb{R}^{n}\to\mathbb{R}$, being $\mu$-a.e. differentiable, such that $\nabla\phi\#\mu=\nu$ solves the Monge problem (4.1). In general, $\phi$ is not smooth and regularity of the optimal transport map is typically studied through the Monge-Ampère equation, e.g., see [ref:philippis2013regularity].
To illustrate what can already be said, we do a simple example.
@@
Suppose we have two zero-mean Gaussian measures $\mu_{0}$ and $\mu_{1}$ on $\mathbb{R}^{n}$, with covariance matrices $\Sigma_{0}$ and $\Sigma_{1}$, respectively. In this case, the optimal transportation map, in the sense of (4.1), is simply $x\mapsto T(x):=Ax$, for $A:=\Sigma_{0}^{-1/2}(\Sigma_{0}^{1/2}\Sigma_{1}\Sigma_{0}^{1/2})^{1/2}\Sigma_{0}^{-1/2}$, that is, $T\#\mu_{0}=\mu_{1}$, e.g., see [ref:peyre2019computational, Rem. 2.31]. In fact, we can construct the interpolation $T(\cdot;s):=(1-s)\mathrm{id}_{\mathbb{R}^{n}}+sT$, with $s\in[0,1]$ and show that for $\mu(\cdot;s):=T(\cdot;s)\#\mu_{0}$ we have that
$\Sigma(s)=\Sigma_{0}^{-1/2}\left((1-s)\Sigma_{0}+s(\Sigma_{0}^{1/2}\Sigma_{1}\Sigma_{0}^{1/2})^{1/2}\right)^{2}\Sigma_{0}^{-1/2}\succ 0.$
Indeed, Brenier’s map is simply $x\mapsto\phi(x;s)=(1-s)\frac{1}{2}\langle x,x\rangle+s\langle x,Ax\rangle$. Now consider the density $\rho(\cdot;s)$ defined through
$\mathbb{R}^{n}\ni x\mapsto\rho(x;s)=\frac{1}{\sqrt{(2\pi)^{n}\mathrm{det}(\Sigma(s))}}e^{-\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle}$
First, note that $s\mapsto\Sigma(s)^{-1}$ is smooth (as captured by the Lie group structure of real invertible matrices, e.g., see [ref:Duistermaat]). Then, set $V(\cdot;s):=-\log\rho(\cdot;s)$. Now suppose that $\Sigma_{1}=I_{n}$. It follows that the family of ODEs $\dot{x}=-\nabla V(x;s)$, as parametrized by $s\in[0,1]$, comprises a homotopy from $\dot{x}=-\nabla V(x;0)$ to $\dot{x}=-x$, through vector fields such that 00 is GAS (e.g., consider the Lyapunov function $x\mapsto\widetilde{V}(x;s):=\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle$). $\circ$
Indeed, the result from Example (4.1) is already known in that sense that for vector fields equipped with convex Lyapunov functions—in particular, linear vector fields—we can construct a stability-preserving homotopy on the level of vector fields [ref:JongeneelSchwan2024TAC]. Note that the fact that we could exploit convexity fits in precisely with work by Caffarelli, e.g., see [ref:caffarelli2000monotonicity, Thm. 11].
@@ 5   Conclusion and future work
We have provided a first step towards addressing Conley’s converse question in generality far beyond our previous work [ref:JongeneelSchwan2024TAC], yet, many open questions remain. Although directly working with flows has its benefits, e.g., see [ref:aguiar2023universal], the main open problem is evidently the extension to vector fields and generic attractors. Several other open questions are as follows.
Open question 1: prove or disprove that (2.1) preserves stability throughout the homotopy.
Open question 2: prove or disprove that Proposition 3.2 holds for $n=5$. Note that a counterexample would disprove the generalized Poincaré conjecture in $\mathsf{Diff}$ for $4$-dimensional spheres.
Although we work with semiflows, we rely on vector fields generating them (see Remark 3.5). Removing this condition is non-trivial, illustrated by [ref:fathi2019smoothing, Sec. 7], but also not futile, as illustrated by the simple example below.
@@
Open question 3: prove or disprove that Proposition 3.2 holds for all semiflows that render 00 GAS.
We also consider relaxing Assumption 3.1, which is instrumental at the moment to be able to appeal to a converse, smooth Lyapunov theory. However, Example 1.4 already showed that the vector field need not be bounded at 00 for a semiflow to exist.
Open question 4: relax Assumption 3.1 as far as possible.
Section 4 just touched upon connections with optimal transport. Motivated by similar work in the context of geometry processing [ref:solomon2019optimalV] (vector field interpolation), we believe this direction deserves a further study.
Open question 5: elucidate what optimal transport can tell us about the existence of stability preserving homotopies on the level of vector fields, plus, what can we tell optimal transport?
At last we point out that classical results typically tell us that certain maps, equivalence classes and so forth, exist. These results, however, rarely provide these objects explicitly. Recent work aims at finding algorithmic schemes to obtain these objects, may it be approximately, e.g., see [ref:bramburger2021deep]. Our work might also benefit from more explicit results, especially when applied in the context of optimization.

### v2

@@ 3   Stability preserving homotopies
Remark 3.8 (Weaker notions of stability).
In general, one cannot relax GAS to mere stability. A reason being that the (Hopf) index of Lyapunov stable equilibria is not fixed [ref:krasnosel1984geometrical, Sec. 52], yet, this index is a homotopy invariant (e.g., for GAS the index is fixed). In Section 4, we do elaborate on ISS. $\circ$
4   Input-to-State Stability
As in [ref:grune1999asymptotic], we can also study “disturbed” systems of the form
$\dot{x}=f(x,d),$ (4.1)
where $f:\mathbb{R}^{n}\times D\to\mathbb{R}^{n}$, with $D\subseteq\mathbb{R}^{m}$, is continuous, and locally Lipschitz on $\mathbb{R}^{n}\setminus\{0\}\times D$. We let $\mathscr{D}_{I}$ denote the set of measurable, locally essentially bounded functions $\mathbb{R}\supseteq I\ni t\mapsto d(t)\in D$, with $d(\cdot)\in\mathscr{D}$ overloading notation, indicating any function of appropriate length. Whenever relevant, we do assume that our solutions are forward complete (e.g., we can appeal to [ref:angeli1999forward, Cor. 2.11]), we denote them using semiflow notation as $\varphi^{t}(x_{0},d(\cdot))$.
Leveraging intuition from linear systems theory, input-to-state stability was derived as a stability notion invariant under coordinate transformations (homeomorphisms that fix 00) [ref:sontagbasicconcepts2008input]. In particular, a system (4.1) is said to be Input-to-State Stable (ISS) when there are $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ such that
$\|\varphi^{t}(x_{0},d(\cdot))\|\leq\max\left\{\beta(\|x_{0}\|,t),\gamma(\|d\|_{\infty})\right\}$
for all $t\geq 0$, $x_{0}\in\mathbb{R}^{n}$ and any $d(\cdot)\in\mathscr{D}$. For the corresponding Lyapunov theory, see [ref:sontag1995characterizations]. Constraining the transient, a system (4.1) is said to be Input-to-State Exponentially Stable (ISES) when there are constants $M,a\in\mathbb{R}_{>0}$ and a $\alpha\in\mathcal{K}_{\infty}$ such that
$\|\varphi^{t}(x_{0},d(\cdot))\|\leq\max\left\{Me^{-at}\|x_{0}\|,\alpha(\|d\|_{\infty})\right\}$ (4.2)
for all $t\geq 0$, $x_{0}\in\mathbb{R}^{n}$ and any $d(\cdot)\in\mathscr{D}$. It can be shown [ref:grune1999asymptotic, Thm. 3], that there is always a coordinate transformation that brings an ISS system into one that is ISES with normalized constants, that is, $M=a=1$.
Inspired by the above, it seems plausible that $\dot{x}=-x+d$ fulfils the role of “canonical ISS system” on $\mathbb{R}^{n}$. For instance, recall that the solution satisfies $\varphi^{t}(x_{0},d(\cdot))=e^{-t}x_{0}+\int^{t}_{0}e^{-(t-\tau)}d(\tau)\mathrm{d}\tau$ for any $t\geq 0$, $x_{0}\in\mathbb{R}^{n}$ and $d(\cdot)\in\mathscr{D}$. Therefore, $\|\varphi^{t}(x_{0},d(\cdot))\|\leq\max\{2|e^{-t}|\|x_{0}\|+2\int^{\infty}_{0}|e^{-(t-\tau)}|\mathrm{d}\tau\|d\|_{\infty}\}$. In that sense, $\dot{x}=-x+d$ is canonically ISS due to the simple comparison functions we can employ. Moreover, set $x\mapsto S(x):=\frac{1}{2}\|x\|_{2}^{2}$, it follows that $\langle\partial_{x}S(x),-x+d\rangle\leq\tfrac{1}{2}(\|d\|_{2}^{2}-\|x\|_{2}^{2})$ for all $x,d\in\mathbb{R}^{n}$ and thus
$\int^{t}_{0}\|\varphi^{\tau}(x_{0},d(\cdot))\|_{2}^{2}\mathrm{d}\tau\leq\|x_{0}\|_{2}^{2}+\int^{t}_{0}\|d(\tau)\|_{2}^{2}\mathrm{d}\tau$ (4.3)
for all $t\geq 0$, $x_{0}\in\mathbb{R}^{n}$ and any $d(\cdot)\in\mathscr{D}$, e.g., see [ref:AvdSL2, Ch. 8]. However, this means that $\dot{x}=-x+d$ can be understood, in particular, as a canonical system with a finite (unitary) linear $L_{2}$ gain. Indeed, [ref:grune1999asymptotic, Thm. 4] shows that an ISES system can be transformed into a system with such an $L_{2}$ gain.
Here, we focus on homotopies through coordinate transformations of $x$ and $d$, for otherwise we could “remove” the disturbance, e.g., consider $f(x,\theta d)$ with $\theta\in[0,1]$. Now, to exemplify why the existence of a homotopy from any ISS system, through coordinate transformation, to $\dot{x}=-x+d$ is too strong, consider the following.
Example 4.1 (No canonical system).
Consider the linear ISS system
$\begin{pmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{pmatrix}=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}\begin{pmatrix}{x}_{1}\\ {x}_{2}\end{pmatrix}+\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\begin{pmatrix}{d}_{1}\\ {d}_{2}\end{pmatrix},$ (4.4)
or in short: $\dot{x}=Ax+R(d)$. As $R\notin\mathrm{Homeo}^{+}(\mathbb{R}^{2};\mathbb{R}^{2})$, there is no continuous path from $\mathrm{id}_{\mathbb{R}^{2}}$ to $R$ (or to $R^{-1}$ for that matter), through coordinate transformations. However, if we pick $x\mapsto S(x):=\tfrac{1}{2}\|x\|_{2}^{2}$ as our storage function then we again find that the system (4.4) satisfies $\langle\partial_{x}S(x),Ax+R(d)\rangle\leq\tfrac{1}{2}(\|d\|_{2}^{2}-\|x\|_{2}^{2})$ and thus, (4.4) satisfies the canonical $L_{2}$ gain bound (4.3). $\circ$
The way to interpret the upcoming proposition is as follows. Given a solution under (4.1), assumed to be ISS. This solution is understood as fixed data and now someone gradually applies a change of coordinates to both $x$ and $d$. It turns out that by doing so, one can always transform this solution data into data satisfying (4.3), thus, into a system like $\dot{x}=-x+d$.
Proposition 4.2 (Homotopy from ISS into a canonical $L_{2}$ gain).
Let $n\neq 5$ and suppose that the system (4.1) is ISS, then we can always find a continuous path of coordinate transformations of both $x$ and $d$, transforming the solutions to (4.1) into solutions of a system satisfying the canonical bound (4.3).
Proof.
First, by [ref:grune1999asymptotic, Thm. 3] and Proposition 3.2, if (4.1) is ISS, we can always find a continuous path of coordinate transformations of $x$, transforming the solutions to (4.1) into solutions of a system that is ISES with normalized constants and admits a representation of the form (4.1).
Second, as we may suppose that (4.1) is ISES with normalized constants, then by [ref:grune1999asymptotic, Thm. 4] and arguments akin to Proposition 3.2, we can always find a continuous path of coordinate transformations of $d$, transforming the solutions to (4.1) into solutions satisfying the canonical $L_{2}$ bound (4.3). To be somewhat explicit, since (4.1) is ISES, then by (4.2), if
$\|x\|_{2}\geq e^{s}\alpha\left(\textstyle\sup_{\tau\in[0,s]}\|d(\tau)\|_{2}\right)$ (4.5)
then $\|\varphi^{t}(x,d(\cdot))\|_{2}\leq e^{-t}\|x\|_{2}$ for all $t\in[0,s]$. Let $V:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$ be defined through $x\mapsto V(x):=\|x\|_{2}^{2}$, then if (4.5) holds we readily get
$V(\varphi^{t}(x,d(\cdot)))\leq e^{-2t}V(x)$ (4.6)
for all $t\in[0,s]$. This must be true in particular for constant disturbances, say $t\mapsto d(t):=d^{\prime}\in D$ for all $t\geq 0$, thus, by subtracting $V(\varphi^{0}(x,d^{\prime}))=e^{-2\cdot 0}\|x\|_{2}^{2}=V(x)$ on both sides of (4.6), dividing by $t$ and considering $t\to 0^{+}$ we get that $\|x\|_{2}\geq\alpha(\|d^{\prime}\|_{2})$ implies that $L_{f(\cdot,d^{\prime})}V(x):=\langle\partial_{x}V(x),f(x,d^{\prime})\rangle\leq-2V(x)\leq-V(x)$. Now, akin to [ref:sontag1995characterizations, Rem. 2.4], define
$\widetilde{\alpha}(r):=\sup_{\|\xi\|_{2}\leq\alpha(r)\,\|u\|_{2}\leq r}2\cdot\langle f(\xi,u),\xi\rangle+V(\xi).$
Evidently, as $\widetilde{\alpha}\in\mathcal{K}_{\infty}$, then for $\|x\|_{2}\geq\alpha(\|d^{\prime}\|_{2})$, we have that
$L_{f(\cdot,d^{\prime})}V(x)\leq-V(x)+\widetilde{\alpha}(\|d^{\prime}\|_{2}).$ (4.7)
On the other hand, for $\|x\|_{2}\leq\alpha(\|d^{\prime}\|_{2})$, (4.7) also holds true. As $d^{\prime}\in D$ was arbitrary, (4.7) holds true for any $d^{\prime}\in D$. Now, let $\widehat{f}(\cdot,v):=f(\cdot,R^{-1}(v))$ and see that if the goal is to get $L_{\widehat{f}(\cdot,v)}V(x)\leq-V(x)+\|v\|_{2}^{2}$, then we could select $\mathbb{R}^{m}\ni z\mapsto R(z):=(\widetilde{\alpha}(\|z\|_{2}))^{1/2}(z/\|z\|_{2})$, with $R(0):=0$, which is $C^{\infty}$ on $\mathbb{R}^{m}\setminus\{0\}$ and continuous at 00 as $\widetilde{\alpha}\in\mathcal{K}_{\infty}$. For $R$ to be orientation preserving, we can again introduce a map $\rho$, e.g., $z\mapsto(\widetilde{\alpha}(\|z\|_{2}))^{1/2}\rho\left(z/\|z\|_{2}\right).$
@@ 5   A view from optimal transport
Suppose there is a smooth Lyapunov function $V$, asserting that $0\in\mathbb{R}^{n}$ is GAS under some ODE $\dot{x}=X(x)$. To refine our understanding beyond Theorem 3.2 and towards a similar result for vector fields, we note that our desire relates to optimal transport (OT). Let $\mu_{0},\mu_{1}\in\mathscr{P}(\mathbb{R}^{n})$ be Borel probability measures on $\mathbb{R}^{n}$. Now suppose that $\mathrm{d}\mu_{0}=\kappa_{0}e^{-V}\mathrm{d}\lambda^{n}$ and $\mathrm{d}\mu_{1}=\kappa_{1}e^{-V_{q}}\mathrm{d}\lambda^{n}$, for $\kappa_{0},\kappa_{1}$ normalization constants. Can we transport $\mu_{0}$ to $\mu_{1}$ along sufficiently regular measures, preserving unimodality? Suppose the answer is yes, then we can construct a path of densities $[0,1]\ni s\mapsto f(x;s)$, with $f(x;0)=e^{-V(x)}$ and $f(x;1)=e^{-V_{q}(x)}$. Now, if $f$ is sufficiently regular, we can conclude that $\nabla\log f(x;s)$ renders 00 GAS, for each fixed $s$. It is particularly interesting that the standard Gaussian measure results in the canonical ODE $\dot{x}=-x$. In what follows we touch upon this viewpoint.
The quadratic “Monge formulation” of OT on $\mathbb{R}^{n}$ is as follows, let $\mu,\nu\in\mathscr{P}_{2}(\mathbb{R}^{n}):=\{\mu\in\mathscr{P}(\mathbb{R}^{n}):\int\|x\|_{2}^{2}\mathrm{d}\mu(x)<+\infty\}$, then we want to solve
$\inf_{T:T\#\mu=\nu}\int_{\mathbb{R}^{n}}\|x-T(x)\|_{2}^{2}\mathrm{d}\mu(x),$ (5.1)
where $T\#\mu$ denotes the pushforward of $\mu$. Brenier [ref:brenier1991polar] showed that if $\mu\ll\lambda^{n}$, then, there is a convex map $\phi:\mathbb{R}^{n}\to\mathbb{R}$, being $\mu$-a.e. differentiable, such that $\nabla\phi\#\mu=\nu$ solves (5.1). We do a simple example.
@@
Given two zero-mean Gaussian measures $\mu_{0}$ and $\mu_{1}$ on $\mathbb{R}^{n}$, with covariance matrices $\Sigma_{0}$ and $\Sigma_{1}$, the optimal map in the sense of (5.1) is simply $x\mapsto T^{\star}(x):=Ax$, for $A:=\Sigma_{0}^{-1/2}(\Sigma_{0}^{1/2}\Sigma_{1}\Sigma_{0}^{1/2})^{1/2}\Sigma_{0}^{-1/2}$, e.g., see [ref:peyre2019computational, Rem. 2.31]. Then, for $\mu(\cdot;s):=T(\cdot;s)\#\mu_{0}$ with $T(\cdot;s):=(1-s)\mathrm{id}_{\mathbb{R}^{n}}+sT$ and $s\in[0,1]$ we have that
$s\mapsto\Sigma(s):=\Sigma_{0}^{-1/2}\left((1-s)\Sigma_{0}+s(\Sigma_{0}^{1/2}\Sigma_{1}\Sigma_{0}^{1/2})^{1/2}\right)^{2}\Sigma_{0}^{-1/2}\succ 0.$
Now consider the density $\mathbb{R}^{n}\ni x\mapsto\rho(x;s):=\kappa(s)^{-1}e^{-\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle}$ for $\kappa(s)^{2}:=(2\pi)^{n}\mathrm{det}(\Sigma(s))$ and set $V(\cdot;s):=-\log\rho(\cdot;s)$. Let $\Sigma_{1}=I_{n}$, then it follows that the family of ODEs $\dot{x}=-\nabla V(x;s)$, comprises a homotopy from $\dot{x}=-\nabla V(x;0)$ to $\dot{x}=-x$, through vector fields such that 00 is GAS (e.g., consider the Lyapunov function $x\mapsto\widetilde{V}(x;s):=\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle$). $\circ$
@@ 6   Conclusion and future work
We have provided a step towards addressing Conley’s converse question in some generality, yet, many open problems remain. Although directly working with flows has benefits, e.g., see [ref:aguiar2023universal], the main open problem is the extension to vector fields and generic attractors. Several other questions are as follows.
Open problem 1: characterize stability of (2.1) throughout the homotopy.
Open problem 2: prove or disprove that Proposition 3.2 holds for $n=5$. A counterexample would disprove the smooth Poincaré conjecture for $\mathbb{S}^{4}$.
Although we work with semiflows, to leverage converse Lyapunov theory we rely on vector fields generating them. Removing this condition is non-trivial, illustrated by [ref:fathi2019smoothing, Sec. 7], but also not futile, as illustrated below.
@@
Open problem 3: prove or disprove that Proposition 3.2 holds for all semiflows that render 00 GAS (not just the ones generated by vector fields).
Section 5 touched upon connections with OT. Motivated by work in the context of geometry processing [ref:solomon2019optimalV], we believe that more work is warranted.
Open problem 4: elucidate what OT can tell us about the existence of stability preserving homotopies on the level of vector fields, and vice versa.
Our work also benefits from more explicit results, e.g., see [ref:bramburger2021deep].